FOX News conducted an online poll of 300 adults in February 2015. One question was “Do you agree there should be more gun control laws?”. 120 adults answered Yes.
In a similar online poll conducted in 2012, 35% of adults supported more gun control laws. In order to determine if the proportion of adults who support more gun control laws in 2015 is higher than 3 years ago, we would like to perform the following hypothesis test:
H_0: p=0.35
H_a: p>0.35
(a) Find the test statistic. (Show work and round the answer to two decimal places)



(b) Determine the P-value for this test. (Show work and round the answer to three decimal places. If you use technology to find the P-value, you have to describe the steps)



(c) Is there sufficient evidence to justify the rejection of at the α=0.05 level? Explain.
Please help me with this!

Answer by ewatrrr(24785)

Ho: p=0.35
Ha: p>0.35 claim
sample proportion = p-hat = 120/300 = 0.4
---
test stat:: z(0.4) = (.4-.35)/sqrt[0.4*0.6/300] = 1.8
-----
p-value = P(z > 1.8) = 0.0359
---------------------
.0359< .05
Conclusion:
Since the p-value is less than 5%, reject Ho.
Evidence supports the claim that the proportion of adults who support more gun
control laws in 2015 is higher than 3 years ago